News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Wednesday 15 September 1999 9.00am EST

During the last 24 hours Chandra passed through the radiation belts with ACIS at the focal point. The next load containing the coordinated observation of HR1099 was uplinked and began operating upon exit from the belts at 2.30PM 9/14. During the communication pass after belt exit it was noted that the High Energy Transmission Grating had not been inserted by the daily load and the cause was traced to two software flags that required re-setting at the same time as the MCE had been turned off manually on 9/13. Following assessment, the grating was re-inserted prior to the start of coordinated observation with Hubble which began as scheduled. This observation will continue for the next 2 days.

Assessment of the ACIS energy resolution degradation continued and measurements with ACIS of the calibration source are now planned on either side of the radiation belts for future passages. A test to operate the CCD's at the intermediate temperature of -60 deg C is under discussion.

Preparations are near completion for the start of the first eclipse season on 9/17. A software patch to set power constants in the flight software to appropriate values was uplinked on 9/14.

Return to Status Reports